1.Kamau completed a job in 45 hours for which 60 hours were allowed. His wage rate per hour is shs 16.

Calculate kamau’s total wage according to :-

a) Halsey scheme                   (2 marks)

b) Halsey-Weir scheme         ( 2 marks)

c) Rowan scheme                   ( 2 marks)


1
Expert's answer
2021-09-24T22:17:55-0400

Solution:

a.). Halsey scheme:

Formula = Hours worked "\times" Rate per hour + (50"\% \times" time saved "\times" rate per hour)

= 45 "\times" 16 + ((50"\% \times" (60 – 45) "\times"16) = 720 + 120 = shs.840

 

b.). Halsey-Weir scheme:

Formula = Formula = Hours worked "\times" Rate per hour + ("33\frac{1}{3}""\% \times" time saved "\times" rate per hour)

= 45 "\times"16 + (("33\frac{1}{3}""\% \times" (60 – 45) "\times"16) = 720 + 79.99 = shs.799.99

 

c.). Rowan scheme:

Formula = Hours worked "\times" Rate per hour + ("\frac{Time\; taken}{Time \;allowed}" "\times" time saved "\times" rate per hrs)


= 45 "\times"16 + ("\frac{45}{60}" "\times"15 "\times"16) = 720 + 180 = shs.900